OpenStudy (anonymous):

Make the appropriate substitution so that \[\int_a^b f(x)~dx=\int_{-1}^1 g(t)~dt\] where \(a,b\) are distinct and arbitrary real numbers in the domain of \(f\). Good practice in calculus/algebra.

OpenStudy (anonymous):

Assume \(f\) is a function such that integrating over these intervals works, etc...

OpenStudy (fibonaccichick666):

there are infinite options...?

OpenStudy (anonymous):

I suppose there are... make the simplest one?

OpenStudy (anonymous):

make a bijection of f: v -> w for all integers being part of bipartite sets

OpenStudy (fibonaccichick666):

let x=t+1 c=a-1 d=b-1

OpenStudy (fibonaccichick666):

yes?

OpenStudy (anonymous):

I seem to have made the mistake of trying to generalize a neat substitution method... Hold on, I'll edit the question.

OpenStudy (anonymous):

I suppose we want \(t = g(x)\) where \(g(b) = d\) and \(g(a) = c\). Are you sure this is always possible?

OpenStudy (anonymous):

Regarding the edit: Essentially, you're asked to transform the integral such that you can integrate over an arbitrary interval in one domain, while keeping the interval fixed in another. Does that make sense?

OpenStudy (anonymous):

Let \(f(x)=1\), then clearly it isn't always the case that \(b-a=d-c\)

OpenStudy (anonymous):

Ever time, I mean

OpenStudy (fibonaccichick666):

so essentially we are mapping f(x) into [-1,1], isn't that a hilberts hotel problem?

OpenStudy (fibonaccichick666):

it can only be done if the set is countable

OpenStudy (anonymous):

Do you mean \[\int_a^b f(x)~dx=\int_{-1}^1 g(t)~dt\] this?

OpenStudy (fibonaccichick666):

http://en.wikipedia.org/wiki/Hilbert%27s_paradox_of_the_Grand_Hotel you sure this doesn't apply?

OpenStudy (dan815):

oh

OpenStudy (anonymous):

The only way we can make a substitution with the function staying the same, that I can readily think of, would be \(t=x+c\)

OpenStudy (anonymous):

Yes @FibonacciChick666 the approach I have in mind has nothing to do with it.

OpenStudy (fibonaccichick666):

hmm ok, but it is almost like a havyside function I guess, but that has specific scenarios

OpenStudy (anonymous):

That would shift the bounds, it wouldn't be able to change the scale of them.

OpenStudy (fibonaccichick666):

heavyside*

OpenStudy (dan815):